Find the value of n that makes the expression divisible by 2,4,and 9
Answer by stanbon(48568)

Find the value of n that makes the expression divisible by 2,4,and 9
---
Certainly 2*4*9= 72 is divisible by 2,4, and 9.
---
You probably want the smallest n divisible by 2,4, and 9.
---
Write the prime-factor form of 2,4, and 9
2 = 2
4 = 2^2
9 = 3^2
-----------------
The least common multiple is 2^2*3^2 = 36
-------------------
Cheers,
Stan H.
PS: Any reply should be sent to stanbon@comcast.net.
The algebra.com feedback program is not working.